OpenStudy (anonymous):

What value of k solves the equation? k^-5=1/32

OpenStudy (anonymous):

\[\frac{1}{k^5}=\frac{1}{32}\iff k^5=32\]

OpenStudy (anonymous):

think of a number that, when raised to the power of 5, gives 32

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

if it not obvious, make a guess yes \(k=2\) because \(2^5=32\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw